Accelerator Update for March 16, 2009 - March 18, 2009 |
Monday March 16 The day shift began with Operations monitoring the Tevatron (TeV) store 6901 (226.19E30) and the Recycler stash (65.6E10), with the Antiproton Source stacking (34.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Operations established beam to the Meson MTest experiment at 9:41 AM. Controls experts reported at 1:34 PM that they had installed a new Pbar PBCOOL front end (FE). They also installed a new PBCOOL VME. At 2:16 PM, Operations reported problems with the new PBCOOL FE: transfers were aborted. Experts investigated. At 2:22 PM, Operations reported problems with the Pbar sequencer. A controls expert investigated. A MI expert began a coalescing check at 2:59 PM. Operations reported at 3:56 PM that the Recycler had stashed antiprotons eight times during the day shift. Operations disabled the MTest beam at 6:02 PM. This ended T988 experiment’s test run for now. TeV experts began an end-of-store crystal collimator study at 7:05 PM. CDF and D0 turned off their high voltage during this study. At 7:38 PM, Operations reported that when they turned off the MTest beam, the AP2 transfer line began suffering from bad beam positions. Operations terminated store 6901 at 8:03 PM. At 8:16 PM, Operations reported that they turned off the Pbar autotune program and restored values saved from the midnight shift. Everything looked good until the autotune program was turned back on and then it again mistuned the AP2 transfer line. Operations put the TeV into shot setup at 8:26 PM. Operations established store 6902 at 9:58 PM with an initial luminosity of 309.28E30. At 11:21 PM, Operations reported that they had rebooted the TeV sector D1 BPM house several times and power cycled it once. This fixed a problem with the orbit stabilization feedback. Operations reported at midnight that the Recycler had stashed six times during the evening shift. Tuesday March 17 The midnight shift began with Operations monitoring store 6902 (204.59E30) and the stash (82.1E10), with Pbar stacking (26.2E10), and with NuMI and MiniBooNE taking beam. At 1 AM, a Linac Klystron RF station (KRF5) tripped off on a LCW flow indication. It reset. At 1:38 AM, a Booster vacuum pump (IPRF24) tripped off and caused BRF15 and 16 to also trip. This happened twice in six minutes. Everything reset. At 2:17 AM, KRF5 tripped off for the first of two times in one hour. It reset. At 3:53 AM, the Booster East Anode Power Supply (EAPS) tripped off because of a BRF13 trip. Operators bypassed the station and brought EAPS back online. At 4:57 AM, Operations reported that a NuMI condensation tank high level alarm had tripped off the NuMI beam permit. Operations contacted experts. Operations prepared MI-65 for access at 6:44 AM. The Neutron Therapy Facility (NTF) reported at 7:52 AM that they were beginning to treat the first of three patients for today at 7:52 AM. Booster RF experts began working on BRF13 at 7:53 AM. At 7:54 AM, KRF7 tripped off. It reset. Operations put Pbar into standby at 7:56 AM. Operations reported at 8:04 AM that the Recycler had stashed six times during the midnight shift. Booster experts brought BRF13 back online at 8:13 AM. At 8:14 AM, BRF19 tripped off. It reset. At 8:30 AM, BRF15 tripped off. It reset. Interlock experts reported at 8:33 AM that they would be testing the manhole safety systems. A NuMI expert reported at 8:37 AM that the condensation tank level was fixed. They came out of access from MI-65. At 9:33 AM, Operations reported that due to bad beam positions in MI, the beam intensity to NuMI had to be lowered. An MI expert tuned. At 9:58 AM, the P1 transfer line beam permit tripped off due to a power supply current drop (V714). Everything reset. At 10 AM, the Linac Buncher tripped off. It reset. At 10:12 AM, BRF5 tripped off. It reset. Operations contacted TeV experts at 10:23 AM about the problems with the sector D1 BPM VME from last night. NTF finished treating patients at 12:13 PM. At 12:23 PM, the NuMI beam permit tripped off due to a power supply trip (E:V100) for the first of two times in eight minutes. Everything reset. Operations informed an expert. At 2:06 PM, Operations reported that the MiniBooNE beam permit started tripping off. Operators lowered the beam intensity and tuned beam. At 2:13 PM, Operations reported that a Linac quadrupole power supply (L:QPSC07) began making noise. A Linac expert investigated. At 2:35 PM, Operations reported bad beam positions in MI and sparking in the Booster RF system. Operators reported that there was Booster damper work going on. Operations terminated store 6902 at 2:41 PM. An instrumentation expert restarted that TeV DCCT FE at 2:45 PM. At 2:46 PM, L:QPS102 tripped off. It reset. Operations put the TeV into shot setup at 2:52 PM. A Recycler expert reported at 3:58 PM that he had rebooted the Recycler dampers. Operations reported at 3:59 PM that the Recycler had stashed six times during the day shift. Operations established store 6903 at 4:22 PM with an initial luminosity of 289.64E30. Operations reported at 6:34 PM that they had received permission to send beam to the new Meson MTest experiment, T971. Water experts reported at 4:35 PM that they had made up water for the LCW systems at MI-60 and 62. At 4:47 PM, Operations reported that FESS experts were investigating LCW pump problems at MI-40 and MI-10. Operations established beam to the MTest T971 experiment at 5:31 PM. Operators disabled the beam about an hour later. A water expert reported at 5:47 PM that the LCW pump at MI-10 would not turn on in remote, but it would run in local. A FESS expert later reported that a cable fault was reported for this pump last year, but was never repaired. Pbar experts began a study at 6:27 PM. At 7:31 PM, MI-60 suffered a vacuum burst. The vacuum valves closed. Operations put Pbar into standby at 7:33 PM. This lasted for about 20 minutes. At 8:33 PM, Operations reported a vacuum burst in Recycler. The emittances grew and antiprotons started to seep away. A recycler expert suggested operators restart the vacuum pumps. Operators turned the pumps on, the pumps faulted off, and the rest of the antiprotons went away. A Recycler expert investigated the vacuum burst and reported at 9:33 PM that this was “a classic example of an argon instability.” It was due, in the past, to a very slow vacuum loss over a long period of time that caused the burst. After the burst the vacuum actually gets better for a while. The last Recycler burst was in August of 2008. Operations reported at 10:36 PM that the Recycler vacuum was good. Operations reported at midnight that the Recycler had stashed three times during the evening shift. Wednesday March 18 The midnight shift began with Operations monitoring store 6903 (114.98E30) and the stash (90.2E10), with Pbar stacking (34.3E10), and with NuMI and MiniBooNE taking beam. At 2:25 AM, the 400 MeV area beam permit tripped off for the first of two times in one minute. It reset. Operators reported that KRF3 was the cause of the trips. Operations halted the MiniBooNE beam. Experts were called in to investigate. Operations reported at 5:52 AM that Linac had resumed sending beam. Unfortunately, the problem cleared itself before the experts could find the cause. At 6:03 AM, LRF4 tripped off. It reset. At 6:14 AM, KRF6 tripped off. It reset. At 6:31 AM, Operations reported that the Recycler damper did something that caused the loss of 20 mA of antiprotons and then returned to stable running. Operations contacted experts. NTF reported at 7:52 AM that they were beginning to treat patients. Operations reported at 8:06 AM that the Recycler had stashed eight times during the midnight shift. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/18/2009 email Fermilab |
| Security, Privacy, Legal | |